4. Calculate [H30*], [OH-], pH and pOH of a 0.5 M CH3COOH solution, Ka = 1.8 x 10-5.
Answer #1

Glution CH3COOH solution is Chls COOH + H2O = CH. C00- + Hoot Equie boium constant (La) = (Chs Coo] [4304] - [ch cool ] In eq1.8x10-5 = x² (0.5-2) as compared too.5 en is very small value co we neglect the n from denominator So 0.5-2 written as 0.5.We know that - Shot] [oH] = 1810-14 [04] = 3x10-14 [43 ot ] [OH = H10th 3x103 = 3.33 X10-12M Tot = 3.33X10 M ARS we know thatDate Page NE PH = 2.523 pH = 2.523 ans. PH+POH = 14 POH = 14 - PH = 14- 20523 =11,477 POH = 11.477 Ang